OpenStudy (help_people):

The function f(t) = t2 + 4t − 14 represents a parabola. Part A: Rewrite the function in vertex form by completing the square. Show your work. Part B: Determine the vertex and indicate whether it is a maximum or a minimum on the graph. How do you know? Part C: Determine the axis of symmetry for f(t).

OpenStudy (help_people):

i have all the parts ecvept for a

OpenStudy (astrophysics):

So you want to complete the square?

OpenStudy (astrophysics):

If you have \[x^2+ax=0\] we can do the following \[x^2+\frac{ 2ax }{ 2 } + \left( \frac{ a }{ 2 } \right)^2-\left( \frac{ a }{ 2 } \right)^2=0\] (complete the square)

OpenStudy (astrophysics):

But don't forget about third term, you just add that, so I'll show what I mean by starting it off  now you can factor